# Source Code Foundation Models are Transferable Binary Analysis Knowledge Bases

### v1

@@ 3.2 Binary Function Name Recovery
We further evaluate the performance of ProRec on the binary function name recovery task. Different from generating summary for a decompiled function, recovering function name requires a tool to have more accurate understanding about program contexts and more concise abstraction for program semantics, as shown by our case study in Appendix E. Please refer to Appendix B.3 for the importance and use scenarios of function name recovery.
@@
We use the source code function names as the ground truth for name recovery. Similar to the binary summarization task, we conduct experiments with three setups: we input to an LLM the decompiled code, the decompiled code with source code snippets obtained by a retriever, and the decompiled code with source code snippets generated by ProRec. The first two setups are considered baselines for comparison.
@@
We use two sets of metrics to evaluate function name. We adapt ROUGE-L [8]for function name since it provides a fine-grained evaluation on the n-gram level. Following existing work in the reverse engineering domain [32], we further evaluate function names by tokenizing pairs of predicted names and ground truth names, and calculating set intersections between the token sets of the predictions and the ground truth. Please see Appendix B.4 for formal definition. For both metrics, we report precision and recall.
@@ 4 Results
In all the following experiments, we report ProRec results based on CodeLlama-34b (4bit quantized). For both the retrieval-augmented baseline (+retrieval) and ProRec (+ProRec), we use their top-5 contexts as augmentation. The versions of the black-box LLMs are gpt-3.5-turbo-1106 for GPT3.5-turbo, claude-3-haiku-20240307 for Claude-3, gemini-1.0-pro for Gemini-Pro, and gpt-4-turbo-2024-04-09 for GPT4 Evaluator.
@@ 4.1 Binary Summarization Results
Figure 4: Scores from GPT4 evaluator for summaries of GPT3.5-turbo. The x-axes denote context relevance (left) and functionality (right), respectively. Larger scores are better. Bars denote the number of summaries with the corresponding score, and dashed lines denote the number of summaries with at least the corresponding score.
We show the results for binary summarization in Table 1. Observe that ProRec helps all models generate better summary in terms of CHRF. A retriever, on the other hand, may introduce noise to a model and even makes the results worse (e.g., for the Gemini-Pro model). Moreover, we can see that ProRec achieves higher scores when evaluated with the GPT4Evaluator on functionality (G4-F) and context relevance (G4-C), indicating the summary of ProRec is more helpful to a human reverse engineer.
@@
On the other hand, we can see that for the context relevance question, both RAG and ProRec introduces more useful context information to the model, and thus the resulting summaries have closer relevance to the groundtruth source code. Especially, queries with the code snippets generated by ProRec achieve more scores 4 and 5 than queries enhanced with a retriever’s results. That illustrates ProRec indeed generates code snippets with better context relevance than a retriever.
For the functionality question, we can observe similar patterns. That indicates better contexts introduced by ProRec help the LLM to understand code functionality. We show a detailed example in Appendix E.
Table 1: Binary summarization results. G4-F and G4-C denote GPT4Evaluator for functionality and context relevance, respectively.
Model CHRF G4-F G4-C
GPT-3.5-turbo - 30.4 3.6 3.8
GPT-3.5-turbo +retrieval 31.7 3.7 3.9
GPT-3.5-turbo +ProRec 33.5 4.2 4.0
Gemini-Pro - 27.1 3.7 3.5
Gemini-Pro +retrieval 26.4 3.4 3.2
Gemini-Pro +ProRec 27.6 3.8 3.6
Claude-3 - 33.5 3.7 3.9
Claude-3 +retrieval 33.9 3.8 3.9
Claude-3 +ProRec 34.9 4.0 4.1
@@ 4.2 Binary Function Name Recovery Results
Table 2: Binary function name recovery results. The following three columns denote the precision, recall, and F1 score proposed by the reverse engineering work SymLM [32]. The last three columns denote the precision, recall, and F1 score for ROUGE-L [8].
SymLM ROUGE-L
Precision Recall F1 Precision Recall F1
GPT-3.5-turbo - 16.3 20.9 17.2 39.8 54.7 44.8
GPT-3.5-turbo +retrieval 15.5 21.3 17.0 39.4 56.5 45.2
GPT-3.5-turbo +ProRec 22.2 28.3 23.5 41.7 58.3 47.3
Gemini-Pro - 25.3 28.7 25.3 44.5 55.3 47.8
Gemini-Pro retrieval 22.7 23.3 21.6 44.2 50.2 45.5
Gemini-Pro +ProRec 32.6 30.5 29.9 50.8 53.8 50.6
Claude-3 - 16.5 22.3 17.9 38.9 48.9 40.3
Claude-3 retrieval 19.9 23.9 20.5 40.6 51.1 43.3
Claude-3 +ProRec 25.9 29.4 26.0 44.6 54.7 47.4
We show results for binary function name recovery in Table 2. We can see that the code snippets generated by ProRec helps all three LLMs predict better names in terms of precision, recall, and F-1 score measured by the metrics designed for reverse engineering task [32]. Especially, ProRec outperforms a retriever by a large margin, indicating that ProRec generates more relevant code than a retriever. For Rouge-L-based metrics, ProRec introduces better performance for GPT-3.5 and Claude-3. However, for Gemini-Pro, the recall is slightly lower than directly query the model without relevant source code snippets. We find that is because Gemini-Pro tends to generate lengthy descriptive names when not provided with specific contexts. And vice versa, the predicted name becomes shorter when provided with specific contexts. For example, without specific context, Gemini-Pro predicts printf_with_variable_arguments for a function printing error messages. After getting ProRec’s results, the prediction becomes error. That means, without context information, Gemini-Pro has higher probability to coincide with the groundtruth tokens since it generates longer names. On the other hand, the precision is significantly lower than ones with better context.
@@ 5 Analysis
5.1 How does black-box LLM’s internal analysis $\mathcal{A}$ help robust recovery?
@@
We study the influence of LLM’s internal analysis by evaluating retrieval-augmented recovery and ProRec with different number of additional contexts, since we believe this kind of internal analysis is crucial for LLM-based recoverers to perform robust binary recovery with additional contexts, especially when the provided contexts are noisy. We run binary function name recovery for 100 randomly sampled examples, 3 times, for zero-shot recovery (decomp-only), retrieval-augmented recovery (retrieval), and ProRec. As shown in Figure 5, the internal analysis consistently reduce the variance of function name recovery performance of both retrieval-augmented recovery and ProRec. This is particularly true for retrieval when $k$ gets large. We deem that it may due to a lack of function-level similar source code in the data store. On the other hand, we observe sometimes LLM tend to be too conservative without leveraging the extra contexts with the internal analysis, potentially because of our not specifically optimized prompts which can be fixed by making some adjustments. Moreover, we argue that misleading is worse than not informative, and reverse engineers can further interact with LLMs for more aggressive induction after obtaining a conservative response.
5.2 Ablation Study on Base Source Code Foundation Model Size
Table 3: Statistics of Prober with Different Base SCFM Sizes
@@ 5.2 Ablation Study on Base Source Code Foundation Model Size
5.3 Correlation Between Human Preference and Auto-Metrics on Binary Summarization
For the binary summarization task, we conduct a human study to measure the correlation between human preference and automatic metrics. We conclude that the automatic metrics CHRF and the GPT4Evaluator are more consistent with human preferences. Please refer to Appendix D for details.
@@ 6 Related Work
6.1 Large Multimodal Models
@@
Recent advancements in vision-language models have demonstrated their efficacy across a range of practical applications such as image captioning [69], visual question answering [5, 16, 3], and image-text matching [40]. While the limited availability of datasets that align different modalities was perceived as a major impediment to scalability, recent works leverage the knowledge embedded within pre-trained large language models, including Flamingo [2], OpenFlamingo [6], Palm-E [19], BLIP-2 [39], along with other existing works [15, 41, 75, 21]. Beyond their capacity to interpret diverse information modalities such as images [62] and audio [27], LLMs have increasingly been aligned with graph structures [11, 58] and gained widespread attention. In particular, there have been successful attempts that leverage LLMs for graph data involves the Graph2Text strategy, which transforms graph data into textual representations. This technique has been effectively utilized in several studies [61, 23, 70]. The designs in the prober of ProRec share some similarity with recent LMMs, with a modality-specific encoder, and a SCFM decoder. However, we tackle the binary-source code multi-modality which is largely unexplored compared to popular modalities. Also, the multi-modal prober is used in a larger probe-and-recover framework instead of end-to-end training.
6.2 Retrieval-Augmented Generation
@@
Retrieval-augmented generation is widely applied in knowledge-intensive scenarios, such as question answering [24, 28, 33], molecule generation [65], and source code generation [74]. By leveraging a non-parametric datastore, retrieval-augmented approaches decompose knowledge and LMs, can complement some long-tail knowledge or keep the knowledge up-to-date without heavy tuning the model which is costly. ProRec, on the other head, try to exploit knowledge within a parametric SCLM for black-box LLM-based binary recovery. A closely related work is GenRead [71] that prompts InstructGPT to generate context instead of retrieval for knowledge intensive tasks. ProRec differs from this work in that binary recovery requires cross-modal understanding and informative contexts cannot be obtained by directly prompting LLMs We introduce specially designed cross-modal alignment to allow informative context generation.
6.3 Binary Reverse Engineering
@@
Advances in machine learning models have been widely used to solve challenging tasks in binary program analysis [51, 49, 60, 57, 50]. However, most work focuses on reasoning binary program, and is not human-oriented. Another stream of work trained smaller end-to-end models for individual human oriented tasks, such as variable name prediction [47, 13] and function name recovery [32]. Nonetheless, these models are not benefiting from pretraining efforts and thus have sub-optimal performance [56]. Preliminary study shows HOBRE remains a challenge for state-of-the-art LLMs [31, 56]. Our efforts attempt to address this challenge, leveraging pre-trained knowledge of binary understanding models and SCFMs to help HOBRE.
@@ 7 Conclusion
We experiment with a simple achitecture and straightforward alignment of the binary-source prober in this paper, which might not be optimal for ProRec. Future work can explore better prober architecture and alignment objectives. Moreover, currently we only focus on intra-procedure analysis, similar to most existing work. In practice, HOBRE needs to deal with full binary with multiple functions. An important direction will be extending ProRec to inter-procedure scenarios, where additional information from the whole program such as call-graph can be leveraged.
@@ Appendix A Details for Training Assembly-Source Code Dual-Encoder
Here, we use cosine similarity for $\text{sim}(\cdot,\cdot)$. In order to have more negative samples, we use momentum encoders [25] for both modality with a queue size 4096. We train the model with learning rate 5e-5, a batch size of 16, 1k warmup steps, and 17k total steps.
@@ B.1 Dataset Quality Assurance and Preprocessing
B.2 Details and Rationale for GPT4Evaluator
@@
We provide GPT4 with the decompiled code, the corresponding source code, the reference summary, and the summary to evaluate. For each question, we adapt the Likert scale  [^5] https://en.wikipedia.org/wiki/Likert_scale and instruct GPT4 to output a score from 1(worst) to 5(best). We give detailed description for each score.
@@
B. Does the summary reflect relevant functionality? Answer the question in range 1(best) to 5(worst). Functionality means the specific high-level behaviors performed in a function (e.g., sort, string comparison, decoding package, printing error messages). For 5, the functionality in the summary should be almost exactly the same to the reference. For 4, the functionalities in the summary are similar to the reference. It may be vague in details, but the overall functionality and purpose is correct. For 3, the summary does not specify functionality. It only repeats some low-level operations without high level abstractions. For 2, the summary specify relevant but inaccurate functionality. The functionality specified in the summary may be relevant to the reference summary, but they have significant differences. For 1, the summary contains irrelevant functionality. It is contains a totally different behavior with the reference. Your output should first briefly comment the summary from the aforementioned perspectives. Do not allow the length of the responses to influence your evaluation. Be as objective as possible.
@@
For other goals in Table 4, goals (1)–(2) are associated with specific analyses, instead of programs. Goal (3) aims to capture the general properties of a program (e.g., the size of a program, the sections in a binary executable file). These properties are easily accessible. The following three goals (4–6) are achieved by a decompiler. The decompiler recovers call to the system APIs (goal 4), reasons instructions and lifts them to a C-like syntax (goal 5), and recovers data dependences by introducing variables (goal 6). Therefore, the focus of binary summarization is on the last two goals, requiring understanding and reasoning of program semantics.

### v2

@@ 3.2 Binary Function Name Recovery
Different from generating summary for a decompiled function, recovering function name requires more accurate understanding about program contexts and more concise abstraction for program semantics. This assists reverse engineers in efficiently navigating numerous functions in a real world binary program. A detailed example is provided in Appendix F.
@@
We use the source code function names as ground truth for name recovery. Similar to the binary summarization task, we conduct experiments with three setups: prompting recoverers with only the decompiled code, with decompiled code and source code snippets obtained by a retriever, with decompiled code and source code snippets generated by ProRec. The first two setups are considered baselines for comparison.
@@
We evaluate the performance of a tool for the binary function name recovery task at different levels of granularity.
Token-level Metrics. In line with existing work in reverse engineering Jin et al. (2022), we tokenize both the predicted function name and the corresponding ground truth, then compute precision, recall, and F1 score at the token level. For each metric, we first calculate the scores for individual function name predictions and then average them across all functions.
Character-level Metrics. We adapt BLEU Papineni et al. (2002), METEOR Banerjee and Lavie (2005), ROUGE-L Lin (2004) for the function name by tokenizing function names into characters and computing these metrics on character level, similar to Lin et al. (2018); Shi et al. (2022). They provide a fine-grained evaluation of the function names and can avoid some limitations of tokenization.
@@ 4 Results
In all the following experiments, we report ProRec results based on CodeLlama-34b (4bit quantized). For both the retrieval-augmented baseline (+retrieval) and ProRec (+ProRec), we use their top-5 contexts as augmentation. The versions of the black-box LLM recoverers are gpt-3.5-turbo-1106 for GPT3.5-turbo, claude-3-haiku-20240307 for Claude-3, gemini-1.0-pro for Gemini-Pro, and gpt-4-turbo-2024-04-09 for GPT4 Evaluator.
@@ 4.1 Binary Summarization Results
[…]
Summarization Function Name Recovery
[…]
GPT-3.5-turbo - 30.4 3.6 3.8 16.3 20.9 17.2 11.9 45.1 38.1
GPT-3.5-turbo +retrieval 31.7 3.7 3.9 15.5 21.3 17.0 10.8 45.5 38.3
GPT-3.5-turbo +ProRec 33.5 4.2 4.0 22.2 28.3 23.5 14.4 47.6 41.1
Gemini-Pro - 27.1 3.7 3.5 25.3 28.7 25.3 16.8 48.1 39.5
Gemini-Pro +retrieval 26.4 3.4 3.2 22.7 23.3 21.6 16.2 45.8 36.5
Gemini-Pro +ProRec 27.6 3.8 3.6 32.6 30.5 29.9 22.4 50.9 40.9
Claude-3 - 33.5 3.7 3.9 16.5 22.3 17.9 13.2 40.7 33.9
Claude-3 +retrieval 33.9 3.8 3.9 19.9 23.9 20.5 14.0 43.3 36.0
Claude-3 +ProRec 34.9 4.0 4.1 25.9 29.4 26.0 17.8 47.7 40.1
We show the results for binary summarization in Table 1. Observe that ProRec helps all models generate better summary in terms of CHRF. A retriever, on the other hand, may introduce noise (irrelevant source functions) to a model and even makes the results worse (e.g., for the Gemini-Pro model). Moreover, we can see that ProRec achieves higher scores when evaluated with the GPT4Evaluator on functionality (G4-F) and context relevance (G4-C), indicating the summary of ProRec is more helpful to a human reverse engineer.
Figure 4: Scores from our proposed GPT4 evaluator for summaries generated basd on GPT3.5-turbo. The x-axes denote context relevance (left) and functionality (right), respectively. Larger scores are better. Bars denote the number of summaries with the corresponding score, and dashed lines denote the number of summaries with at least the corresponding score.
@@
Table 2: Human evaluation of binary summarization results w.r.t. context relevance and functionality.
Cont. Rel. Func.
- 4.29 4.22
 +retrieval 4.49 4.43
 +ProRec 4.76 4.62
On the other hand, we can see that for the context relevance question, both retrieval-augmented baseline and ProRec introduces more useful context information to the model, and thus the resulting summaries have closer relevance to the ground truth source code. Especially, queries with the code snippets generated by ProRec achieve more scores 4 and 5 than queries enhanced with a retriever’s results. That illustrates ProRec indeed generates code snippets with better context relevance than a retriever.
For the functionality question, we can observe similar patterns. That indicates better contexts introduced by ProRec help the LLM to understand code functionality. We show a detailed example in Appendix F.
Human Evaluation
Table 2 presents the human evaluation results from our user study for 50 randomly sampled summaries of each method (using GPT-3.5-turbo as the recoverer). The human evaluation aligns with the automatic metrics: ProRec consistently outperforms the other approaches in terms of both context relevance and functionality, according to human judgment.
@@ 4.2 Binary Function Name Recovery Results
We show results for binary function name recovery in Table 1. We can see that the code snippets generated by ProRec helps all three LLMs predict better names in terms of token-level precision, recall, and F-1 score designed for reverse engineering task Jin et al. (2022). Especially, ProRec outperforms a retriever by a large margin, indicating that ProRec generates more relevant code than a retriever. For character-level metrics, ProRec shows similar improvements, not biasing towards certain metrics. Note that different LLM recoverers can have different performance on the two tasks, e.g., Gemini-Pro is better at binary function name recovery than summarization compared to other two models, yet this does not influence the improvement ProRec brings to both tasks and all recoverers.
@@ 5 Analysis
How does black-box LLM’s internal analysis $\mathcal{A}$ help robust recovery?
@@
We study the influence of LLM’s internal analysis by evaluating retrieval-augmented recovery and ProRec with different number of additional contexts, since we believe this kind of internal analysis is crucial for LLM-based recoverers to perform robust binary recovery with additional contexts, especially when the provided contexts are noisy. We run binary function name recovery for 100 randomly sampled examples, 3 times, for zero-shot recovery (dec-only), retrieval-augmented recovery (recovery), and ProRec. As shown in Figure 5, the internal analysis consistently reduce the variance of function name recovery performance of both retrieval-augmented recovery and ProRec. This is particularly true for retrieval when $k$ gets large. We deem that it may due to a lack of function-level similar source code in the data store. On the other hand, we observe sometimes LLM tend to be too conservative without leveraging the extra contexts with the internal analysis, potentially because of our not specifically optimized prompts which can be fixed by making some adjustments. Moreover, we argue that misleading is worse than not informative, and reverse engineers can further interact with LLMs for more aggressive induction after obtaining a conservative response.
Ablation Study on Base Source Code Foundation Model Size
Table 3: Statistics of prober with different base SCFM sizes.
@@
Case Study
We examine three specific cases to illustrate the performance and limitation of ProRec in Appendix F.
@@ 6 Related Work
Large Multimodal Models
@@
Recent advancements in vision-language models have demonstrated their efficacy across a range of practical applications such as image captioning Xu et al. (2015), visual question answering Antol et al. (2015); Das et al. (2017); Anderson et al. (2018), and image-text matching Li et al. (2019). While the limited availability of datasets that align different modalities was perceived as a major impediment to scalability, recent works leverage the knowledge embedded within pre-trained large language models Alayrac et al. (2022); Awadalla et al. (2023); Li et al. (2023); Dai et al. (2024); Liu et al. (2023); Zhu et al. (2023). Beyond their capacity to interpret diverse information modalities such as images Wang et al. (2024b) and audio Huang et al. (2024), LLMs have increasingly been aligned with graph structures Chai et al. (2023); Tang et al. (2023) and gained widespread attention. In particular, there have been successful attempts that leverage LLMs for graph data involves the Graph2Text strategy, which transforms graph data into textual representations. This technique has been effectively utilized in several studies Wang et al. (2024a); Guo et al. (2023); Ye et al. (2023a). The designs in the prober of ProRec share some similarity with recent LMMs, with a modality-specific encoder, and a SCFM decoder. However, we tackle the binary-source code multi-modality which is largely unexplored compared to popular modalities. Also, the multi-modal prober is used in a larger probe-and-recover framework instead of end-to-end training.
Retrieval-Augmented Generation
@@
Retrieval-augmented generation is widely applied in knowledge-intensive scenarios, such as question answering Guu et al. (2020); Izacard and Grave (2021); Karpukhin et al. (2020), molecule generation Wang et al. (2022b), and source code generation Zhou et al. (2022). By leveraging a non-parametric datastore, retrieval-augmented approaches decompose knowledge and LMs, can complement some long-tail knowledge or keep the knowledge up-to-date without heavy tuning the model which is costly. ProRec, on the other head, tries to exploit knowledge within a parametric SCLM for black-box LLM-based binary recovery. A closely related work is GenRead Yu et al. (2022) that prompts InstructGPT to generate context instead of retrieval for knowledge intensive tasks. ProRec differs from this work in that binary recovery requires cross-modal understanding and informative contexts cannot be obtained by directly prompting LLMs We introduce specially designed cross-modal alignment to allow informative context generation.
Binary Reverse Engineering
@@
Advances in machine learning models have been widely used to solve challenging tasks in binary program analysis Pei et al. (2020, 2021a); Wang et al. (2022a); Xu et al. (2023a); Su et al. (2024); Pei et al. (2021b). However, most work focuses on reasoning binary program, and is not human-oriented. Another stream of work trained smaller end-to-end models for individual human oriented tasks, such as variable name prediction Pal et al. (2024); Chen et al. (2022a); Xu et al. (2023b, c), and function name recovery Jin et al. (2022). Nonetheless, these models are not benefiting from pretraining efforts and thus have sub-optimal performance Shang et al. (2024). Preliminary study shows HOBRE remains a challenge for state-of-the-art LLMs Jin et al. (2023); Shang et al. (2024). Our efforts attempt to address this challenge, leveraging pretraining knowledge of SCFMs to help HOBRE.
@@ 7 Conclusion
We experiment with a simple achitecture and straightforward alignment of the binary-source prober in this paper, which might not be optimal for ProRec. Future work can explore better prober architecture and alignment objectives. Moreover, currently we only focus on intra-procedure analysis, similar to most existing work. In practice, HOBRE needs to deal with full binary with multiple functions. An important direction will be extending ProRec to inter-procedure scenarios, where additional information from the whole program such as call-graph can be leveraged, building program-level binary reverse engineering agents.
8 Acknowledgement
We thank the anonymous reviewers for their valuable comments and suggestions. We thank all the participants in our user study. We are grateful to the Center for AI Safety for providing computational resources. This research was supported in part by DARPA VSPELLS - HR001120S0058, IARPA TrojAI W911NF-19-S-0012, NSF 1901242 and 1910300, ONR N000141712045, N000141410468 and N000141712947. Any opinions, findings, and conclusions in this paper are those of the authors only and do not necessarily reflect the views of our sponsors.
@@ Appendix A Details for Training Assembly-Source Code Dual-Encoder
Here, we use cosine similarity for $\text{sim}(\cdot,\cdot)$. In order to have more negative samples, we use momentum encoders [23] for both modality with a queue size 4096 and momentum 0.999. We train the model with learning rate 5e-5, a batch size of 16, 1k warmup steps, and 17k total steps.
@@ B.1 Dataset Quality Assurance and Preprocessing
B.2 Evaluation Aspects and Rationale for Human Study and GPT4Evaluator
@@
We provide GPT4 with the decompiled code, the corresponding source code, and the reference summary to evaluate. For each question, we adapt the Likert scale  [^6] https://en.wikipedia.org/wiki/Likert_scale and instruct GPT4 to output a score from 1 (worst) to 5 (best). The users in our human study are provided with similar information as questionnaires to perform human judgment of summaries.
@@
B. Does the summary reflect relevant functionality? Answer the question in range 5(best) to 1(worst). Functionality means the specific high-level behaviors performed in a function (e.g., sort, string comparison, decoding package, printing error messages). For 5, the functionality in the summary should be almost exactly the same to the reference. For 4, the functionalities in the summary are similar to the reference. It may be vague in details, but the overall functionality and purpose is correct. For 3, the summary does not specify functionality. It only repeats some low-level operations without high level abstractions. For 2, the summary specify relevant but inaccurate functionality. The functionality specified in the summary may be relevant to the reference summary, but they have significant differences. For 1, the summary contains irrelevant functionality. It is contains a totally different behavior with the reference. Your output should first briefly comment the summary from the aforementioned perspectives. Do not allow the length of the responses to influence your evaluation. Be as objective as possible.
@@
For other goals in Table 4, goals (1–2) are associated with specific analyses, instead of programs. Goal (3) aims to capture the general properties of a program (e.g., the size of a program, the sections in a binary executable file). These properties are easily accessible. The following three goals (4–6) are achieved by a decompiler. The decompiler recovers call to the system APIs (goal 4), reasons instructions and lifts them to a C-like syntax (goal 5), and recovers data dependences by introducing variables (goal 6). Therefore, the focus of binary summarization is on the last two goals, requiring understanding and reasoning of program semantics.